Find the acute angles of the right triangle with sides measuring: 3,4,5. I know the hypotenuse is 5 as it is the longest measurement. I tried putting the other numbers in to get the sine SinA b/5 SinA 4/5 =.0139 but the angle of this is not the right answer. I am having trouble understanding how to use sin, cos, and tan to find answers regarding angles and sides.
Thanks so much for all your time and help.



Answer by stanbon(75887)

Find the acute angles of the right triangle with sides measuring: 3,4,5. I know the hypotenuse is 5 as it is the longest measurement. I tried putting the other numbers in to get the sine SinA b/5 SinA 4/5 =.0139 but the angle of this is not the right answer. I am having trouble understanding how to use sin, cos, and tan to find answers regarding angles and sides.
---
Draw the figure.
Let the hypotenuse be 5.
Let the angle opposite 4 be A.
Then sin-1(4/5) = 53.13 degrees or 0.9273 radians
----
The other acute angle is sin^-1(3/5) = 36.87 degrees or 0.6435 radians
-----
Notice that the angles are complimentary.
===============
Cheers,
Stan H.